# A comprehensive analysis toward the Fermi-LAT source 4FGL J1846.90227 Jets of a proto-planetary nebula producing -rays?

### v1

@@ 3.1 The nature of the uncatalogued radio source
In this subsection, we analyze the possibility that the new radio source is a SNR, for which a radio spectral index study is carried out. We use the averaged image between 1.4 and 1.8 GHz from the THOR survey444Retrieved from https://thorserver.mpia.de/data/continuum/.. This image was convolved with a (FWHM) Gaussian kernel to match the resolution of the GLOSTAR image.
@@ 3.2 The MYSO candidate MSX G30
The MYSO candidate MSX G30 is cataloged as a massive protostar candidate with a bolometric luminosity of about Ā Lā (Lumsden etĀ al., 2013). It is well known that MYSOs have associated massive bipolar molecular outflows which can produce strong shocks when interact with the surrounding medium. In these conditions particle acceleration at relativistic velocities can occur leading to gamma ray emission, as some theoretical models predict (Bosch-Ramon etĀ al., 2010, and references therein).
Crossing the Fermi First Year Catalog with some catalogs of MYSO candidates, Munar-Adrover etĀ al. (2011) obtained a list of MYSOs that are spatially coincident with Fermi sources, and therefore, potentially responsible for the -ray emission. In particular, the authors found a spatial correlation between the Fermi source 1FGL J1846.80233c (a.k.a. 4FGL J1846.90227) and MSX G30. Therefore, we searched in the literature for evidence of molecular outflow activity associated with this MYSO candidate.
Using intermediate angular resolution data (about 15) from the 13CO/C18O (3ā2) Heterodyne Inner Milky Way Plane Survey (CHIMPS; Rigby etĀ al., 2016), Yang etĀ al. (2018) searched for massive outflows associated with a sample of 919 ATLASGAL clumps. In particular, the authors did not find any evidence of outflow activity towards the source AGAL G030.19780.1683, which is the dust condensation in which MSX G30 is embedded. On the other hand, we searched for Extended Green Objects (EGOs) in the catalog of Cyganowski etĀ al. (2008). EGOs are usually associated with massive molecular outflow activity. However, there is not any EGOs cataloged in the region.
In this context, we analyzed high resolution and sensitivity continuum and line ALMA data at Band 6 searching for traces of molecular outflow activity at core scale.
@@
Figureā4 shows a beam average spectrum ranging from about 224.3 to 226.3Ā GHz (Band 6) towards the core mm1. Several molecular lines have been detected in this spectral window, but for the purpose of our analysis we will focus on two of them: CH3OCHO (18,17) and H2CO (3ā2). Figureā5-left panel shows the integrated emission map of CH3OCHO (18,17) E transition. The spatial distribution of this molecule appears well localized onto the position of the core mm1. The mere presence of this transition, whose typical desorption temperature is about 120 K (e.g. Busch etĀ al. 2022), shows that mm1 is a hot molecular core, which confirms that active star formation is ongoing inside.
Figureā5-right panel shows the integrated emission map of H2CO J=3ā2 transition. It can be seen that the spatial distribution of this molecule exhibits a main peak which perfectly match the position of the core mm1. In addition, it can be noticed two emission features with no associated continuum emission, which are perfectly aligned with the core mm1 (see dotted black line in Fig.ā5-right). Taking into account that this molecule usually traces dense gas and molecular outflows (Tychoniec etĀ al., 2021; Okoda etĀ al., 2020), it is very likely that both structures outside the core position are showing signs of outflow activity. In particular, both structures could be tracing the densest gas of the molecular outflows, likely the caps. Given that the velocity difference between both caps are about 3Ā km s-1, we can infer that the lobes are extended mostly at the plane of the sky. The lineal distance between both caps are about 0.5Ā pc, which is in agreement with typical outflows sizes.
@@
We note that 4XMM J184651.2022504 is the XMM-Newton counterpart of the Swift source SwXF4 J184651.6022507 mentioned in Sect.ā1 (see Fig.ā2). According to Kerby etĀ al. (2021), this source has an X-ray flux three orders of magnitude smaller than SwF4 J184650.7022904 (see Sect.ā3.4) and it was characterized as a likely blazar.
This faint X-ray source is related to the mid-IR source WISE J184651.22022505.5 from the All WISE data release Catalog of Cutri etĀ al. (2021). The reported magnitudes for this source in the four mid-IR observed bands are: [3.4]=10.50, [4.6]=10.46, [12.0]=10.08, and [22.0]=5.09 (quality flags AAUB). Even though the quality at the 12 m band indicates that it is an upper limit, it can be used to locate this source in a color-color ([3.4]-[4.6] vs. [4.6]-[12.0]) diagram. This kind of diagrams are very frequently used to determine a blazar association (e.g. DāAbrusco etĀ al. 2012). The source lies very far from the blazars region in such a diagram, in fact, it is located in the region in which the stars are expected to be. Moreover, giving that the source lies very close to the galactic plane (l = 0.08 deg), a galactic origin is more likely. Hence, we discard the blazar nature and its responsibility for the high-energy emission.
@@
Source 4XMM J184650.6022907 is the XMM-Newton counterpart of the Swift source SwXF4 J184650.7022904 (see Fig.ā2). According to the XMM-Newton Serendipitous Source Catalog, Thirteenth Data Release555Catalog available through the XMM-Newton Science Archive interface at https://www.cosmos.esa.int/web/xmm-newton/xsa (4XMM-DR13), this source presents significant variability: the EPIC count-rate between 0.2 and 12 keV increases from for Obs1 (2011) to for observation Obs2 (2018).
Looking for possible counterparts of 4XMM J184650.6022907, we found that it is related to mid-IR source WISE J184650.7-022904 (Cutri etĀ al., 2021) with magnitudes [3.4]=8.861, [4.6]=6.951, [12.0]=7.108, and [22.0]=8.106 (quality flags AAAU). Again, from a color-color ([3.4]-[4.6] vs. [4.6]-[12.0]) diagram, this source is not compatible with a blazar. We also found that the star BD02ā4739 (at 18:46:50.69 02:29:07.30, J2000), cataloged as a spectroscopic binary in the Non-single stars Catalog by GAIA DR3 (Gaia Collaboration, 2022), is considered the optical counterpart in the 4XMM-DR13 catalog. Indeed, it lies at the same coordinate of 4XMM J184650.6022907 within a positional difference less that 0.1 arcsec (see Fig.ā6). Taking into account this perfect positional coincidence, it is very likely that both sources are related, and hence, it deserves more analysis in the context of this work.
@@ 3.4.1 Spectroscopic binary star BD02ā4739
As noted in previous section, it is very likely that the spectroscopic binary star BD02 4739 is the optical counterpart of 4XMM J184650.6022907. The distance and orbital period reported by GAIA DR3 are and , respectively (Gaia Collaboration, 2022). Several other optical/ir associations are reported in the 4XMM-DR13 catalog and Simbad for this star, which are useful to determine its spectral type through a spectral energy distribution (SED) analysis.
In a binary system, in principle both components contribute to the fluxes observed at different wavelengths, hence a SED study must be done with care. In this case, we conducted a SED analysis based on the statement that one of the components of this binary system should be a black hole, a neutron star or a white dwarf, issue discussed in Sect.ā4, and hence we assume that the analyzed fluxes for the SED came mainly from a donor stellar companion.
In Tableā2 we report the associated sources from the WISE Source Catalog (AllWISE, band W4 is excluded because the flux is reported as an upper limit), the Galactic Legacy Infrared Mid-Plane Survey Extraordinaire (GLIMPSE) performed with the Spitzer Space Telescope, the Two Micron All-Sky Survey (2MASS), the GAIA Data Release 3 (Gaia DR3), the Tycho 2 Main Catalog (Tycho-2), and the XMM-Newton Optical Monitor (XMM-OM) Source Catalog (XMM-SUSS5.0). We have excluded the optical source 1237668687974826071 from the Sloan Digital Sky Survey Data Release 12 (SDSS DR12) because several photometric flags are reported. Flux densities () were calculated from the magnitudes reported in the catalogs as , where is the zero-point flux, whose values were obtained from Rodrigo & Solano (2020).
Survey / Source ID | Filter / [m] | [ |
allWISE / | W1 / 3.35 | |
184650.68022907.4 | W2 / 4.60 | |
W3 / 11.56 | ||
Spitzer GLIMPSE / | IRAC 3.6 / 3.54 | |
G030.196500.111 | IRAC 4.5 / 4.48 | |
IRAC 5.8 / 5.70 | ||
IRAC 8.0 / 7.80 | ||
2MASS / | J / 1.24 | |
184650690229072 | H / 1.65 | |
Ks / 2.16 | ||
Gaia DR3 / | BP / 0.51 | |
4259072330161989504 | G / 0.62 | |
RP / 0.77 | ||
Tycho-2 / | B / 0.42 | |
TYC2 5118-160-1 | V / 0.53 | |
XMM-OM / | UVM2 / 0.23 | |
@@
We fit the spectral energy distribution (SED) using a black-body function without interstellar extinction, a reasonable approximation taking into account the proximity of the source. The fit parameters are the effective temperature and the dimensionless normalization :
@@
where is the radius and the distance. In Fig.ā7 we show the fluxes of Tableā2 and the result of the fit. We obtain and (errors correspond to 90% confident level). Using Eq.ā2 we obtain a stellar radius of about .
We estimate the flux by numerically integrating between and and obtain . The bolometric luminosity is estimated from the relation . To determine the spectral type of BD02 4739, we construct the HR diagram of Fig. 8, where we include the location of the main sequence, sub-giant, and giant stars, corresponding to luminosity classes V, IV and, III, respectively (taken from de Jager & Nieuwenhuijzen 1987).
Hence, we suggest that the donor star, this is, the member that would contribute the most to the IR and optical fluxes, should be a low-mass star out of the main sequence, likely a sub-giant star.
@@ 3.4.2 X-ray spectral analysis
We extracted the X-ray spectra of 4XMM J184650.6022907 from the 3 cameras of each XMM-Newton observation, yielding to a set of 6 spectra. The source extraction region was a circle of Ā in radius and background spectra were extracted from nearby circular regions in the same CCD as the source region. We use the evselect task of SAS with a minimum of 15 counts per bin. We obtained 209 (mos1), 245 (mos2), and 416 (pn) counts for observation Obs1, and 359 (mos1), 391 (mos2), and 415 (pn) counts for Obs2, in the energy band. The spectra were displayed and analyzed with the Xspec software (version 12.14.0), which is included in the HEASoft package.
We visually inspected the 6 spectra and found X-ray emission coming from a soft and featureless continuum, without evidence of spectral lines (Fig. 9). We note, however, that the absence of lines in the spectra may be a consequence of the relative low number of counts and the binning. We first fit each observation separately with different emission models in the energy band, to look for possible variations of model parameters between Obs1 and Obs2. To explore the non-thermal scenario we use an absorbed power law (). We obtain an unrealistic soft photon index ( and for Obs1 and Obs2, respectively) and a quite poor fit ( and 1.8, for Obs1 and Obs2)666To characterize the goodness of the fit, we report the reduced , i.e. , where are the degrees of freedom.. To explore the thermal nature of the source, we fitted the spectra with a blackbody, a bremsstrahlung, and an optically thin plasma. We obtained poor results for the blackbody ( and 2.0, for Obs1 and Obs2) and bremsstrahlung ( for Obs1 and Obs2), and an acceptable fit for the thin plasma (). For this model, the 90% confidence range for the hydrogen column densities are (Obs1) and (Obs2), showing no significant variation of the absorbing material between observations. For the electron temperature (), we obtain confidence ranges (Obs1) and (Obs2), implying a slightly warmer plasma for Obs2. Metal abundance () is sub-solar and confidence ranges are for Obs1 and for Obs2.
Based on the results of the individual fit of each observation, and in order to increase the fit statistic, we simultaneously fit Obs1 and Obs2 spectra using a thermal plasma model with the same . In Table 3 we show the fitting results for the absorbed thermal plasma, considering two procedures: i) all parameters except normalization are tie together between Obs1 and Obs2, and ii) only NH is tie together between the observations.
@@
The results of the spectral fitting reported in Table 3 show that the X-ray emission from 4XMM J184650.6022907 is likely produced by a thermal plasma with a temperature and sub-solar abundance. There is a slight increase in the plasma temperature between observations (refer to the right column of the Table), and no significant variations of the metal abundance. The normalization increases by a factor between observations, i.e. in a time scale of 7 years. This indicates that the increase of the flux is produced by an increase of the amount of emitting plasma.
3.5 PN candidate IRAS 18443-0231

### v2

@@ 3.1 The nature of the uncatalogued radio source
In this subsection, we analyse the possibility that the new radio source is a SNR, for which a radio spectral index study is carried out. We use the averaged image between 1.4 and 1.8 GHz from the THOR survey444Retrieved from https://thorserver.mpia.de/data/continuum/.. This image was convolved with a (FWHM) Gaussian kernel to match the resolution of the GLOSTAR image.
@@ 3.2 The MYSO candidate MSX G30
The MYSO candidate MSX G30 is catalogued as a massive protostar candidate with a bolometric luminosity of about Ā Lā (Lumsden etĀ al., 2013). It is well known that MYSOs have associated massive bipolar molecular outflows which can produce strong shocks when they interact with the surrounding medium. In these conditions particle acceleration at relativistic velocities can occur leading to -ray emission, as some theoretical models predict (Bosch-Ramon etĀ al., 2010, and references therein).
Crossing the Fermi First Year Catalogue with some catalogues of MYSO candidates, Munar-Adrover etĀ al. (2011) obtained a list of MYSOs that are spatially coincident with Fermi sources, and therefore, potentially responsible for the -ray emission. In particular, the authors found a spatial correlation between the Fermi source 1FGL J1846.80233c (a.k.a. 4FGL J1846.90227) and MSX G30. Therefore, we searched in the literature for evidence of molecular outflow activity associated with this MYSO candidate.
Using intermediate angular resolution data (about 15) from the 13CO/C18O (3ā2) Heterodyne Inner Milky Way Plane Survey (CHIMPS; Rigby etĀ al., 2016), Yang etĀ al. (2018) searched for massive outflows associated with a sample of 919 ATLASGAL clumps. In particular, the authors did not find any evidence of outflow activity towards the source AGAL G030.19780.1683, which is the dust condensation in which MSX G30 is embedded. On the other hand, we searched for Extended Green Objects (EGOs) in the catalogue of Cyganowski etĀ al. (2008). EGOs are usually associated with massive molecular outflow activity. However, there is not any EGOs catalogued in the region.
In this context, we analysed high resolution and sensitivity continuum and line ALMA data at Band 6 searching for traces of molecular outflow activity at core scale.
@@
Figureā4 shows a beam average spectrum ranging from about 224.3 to 226.3Ā GHz (Band 6) towards the core mm1. Several molecular lines have been detected in this spectral window, but for the purpose of our analysis we will focus on two of them: CH3OCHO (18,17) and H2CO (3ā2). Figureā5-left panel shows the integrated emission map of CH3OCHO (18,17) E transition. The spatial distribution of this molecule appears well localised onto the position of the core mm1. The mere presence of this transition, whose typical desorption temperature is about 120 K (e.g. Busch etĀ al. 2022), shows that mm1 is a hot molecular core, which confirms that active star formation is ongoing inside.
Figureā5-right panel shows the integrated emission map of H2CO J=3ā2 transition. It can be seen that the spatial distribution of this molecule exhibits a main peak which perfectly match the position of the core mm1. In addition, it can be noticed two emission features with no associated continuum emission, which are perfectly aligned with the core mm1 (see dotted black line in Fig.ā5-right). Taking into account that this molecule usually traces dense gas and molecular outflows (Tychoniec etĀ al., 2021; Okoda etĀ al., 2020), it is very likely that both structures outside the core position are showing signs of outflow activity. In particular, both structures could be tracing the densest gas of the molecular outflows, likely the caps. Given that the velocity difference between both caps are about 3Ā kmās-1, we can infer that the lobes are extended mostly at the plane of the sky. The lineal distance between both caps are about 0.5Ā pc, which is in agreement with typical outflows sizes.
@@
We note that 4XMM J184651.2022504 is the XMM-Newton counterpart of the Swift source SwXF4 J184651.6022507 mentioned in Sect.ā1 (see Fig.ā2). According to Kerby etĀ al. (2021), this source has an X-ray flux three orders of magnitude smaller than SwF4 J184650.7022904 (see Sect.ā3.4) and it was characterised as a likely blazar.
This faint X-ray source is related to the mid-IR source WISE J184651.22022505.5 from the AllWISE data release Catalogue of Cutri
etĀ al. (2021). The reported magnitudes for this source in the four mid-IR observed bands are: [3.4]=10.50, [4.6]=10.46, [12.0]=10.08, and [22.0]=5.09 (quality flags AAUB). Even though the quality at the 12 m band indicates that it is an upper limit, it can be used to locate this source in a colour-colour ([3.4]-[4.6] vs. [4.6]-[12.0]) diagram. This kind of diagrams are very frequently used to determine a blazar association (e.g. DāAbrusco etĀ al. 2012). The source lies very far from the blazars region in such a diagram, in fact, it is located in the region in which the stars are expected to be. Moreover, giving that the source lies very close to the galactic plane (l = 0.08 deg), a galactic origin is more likely. Hence, we discard the blazar nature and its responsibility for the high-energy emission.
@@
Source 4XMM J184650.6022907 is the XMM-Newton counterpart of the Swift source SwXF4 J184650.7022904 (see Fig.ā2). According to the XMM-Newton Serendipitous Source Catalogue, Thirteenth Data Release555Catalogue available through the XMM-Newton Science Archive interface at https://www.cosmos.esa.int/web/xmm-newton/xsa (4XMM-DR13), this source presents significant variability: the EPIC count-rate between 0.2 and 12 keV increases from for Obs1 (2011) to for observation Obs2 (2018).
Looking for possible counterparts of 4XMM J184650.6022907, we found that it is related to mid-IR source WISE J184650.7-022904 (Cutri etĀ al., 2021) with magnitudes [3.4]=8.861, [4.6]=6.951, [12.0]=7.108, and [22.0]=8.106 (quality flags AAAU). Again, from a colour-colour ([3.4]-[4.6] vs. [4.6]-[12.0]) diagram, this source is not compatible with a blazar. We also found that the star BD02ā4739 (at 18:46:50.69 02:29:07.30, J2000), catalogued as a spectroscopic binary in the Non-single stars Catalogue by GAIA DR3 (Gaia Collaboration, 2022), is considered the optical counterpart in the 4XMM-DR13 catalogue. Indeed, it lies at the same coordinate of 4XMM J184650.6022907 within a positional difference less that 0.1 arcsec (see Fig.ā6). Taking into account this perfect positional coincidence, it is very likely that both sources are related, and hence, it deserves more analysis in the context of this work.
@@ 3.4.1 Spectroscopic binary star BD02ā4739
As noted in previous section, it is very likely that the spectroscopic binary star BD02 4739 is the optical counterpart of 4XMM J184650.6022907. The distance and orbital period reported by GAIA DR3 are and , respectively (Gaia Collaboration, 2022). Several other optical/ir associations are reported in the 4XMM-DR13 catalogue and Simbad for this star, which are useful to determine its spectral type through a spectral energy distribution (SED) analysis.
In a binary system, in principle both components contribute to the fluxes observed at different wavelengths, hence a SED study must be done with care. In this case, we conducted a SED analysis based on the statement that one of the components of this binary system should be a black hole, a neutron star or a white dwarf, issue discussed in Sect.ā4, and hence we assume that the analysed fluxes for the SED came mainly from a donor stellar companion.
In Tableā2 we report the associated sources from the WISE Source Catalogue (AllWISE, the flux at band W4 is reported as an upper limit), the Galactic Legacy Infrared Mid-Plane Survey Extraordinaire (GLIMPSE) performed with the Spitzer Space Telescope, the Two Micron All-Sky Survey (2MASS), the GAIA Data Release 3 (Gaia DR3), the Tycho 2 Main Catalogue (Tycho-2), and the XMM-Newton Optical Monitor (XMM-OM) Source Catalogue (XMM-SUSS5.0). We have excluded the optical source 1237668687974826071 from the Sloan Digital Sky Survey Data Release 12 (SDSS DR12) because several photometric flags are reported. Flux densities () were calculated from the magnitudes reported in the catalogues as , where is the zero-point flux, whose values were obtained from Rodrigo & Solano (2020). The fluxes were corrected for extinction, which was obtained from the Gaia-2MASS 3D maps of Galactic interstellar dust of Lallement etĀ al. (2019). In the line of sight of BD02ā4739 (, ), we obtained a visual extinction for a distance of 666Extinction calculator available at: https://astro.acri-st.fr/gaia_dev/. corresponds to a reddening using the standard .. In Tableā2 we report the dereddened fluxes (), calculated using the VO SED Analyzer (VOSA777Available at: http://svo2.cab.inta-csic.es/theory/vosa/, Bayo etĀ al. 2008).
Survey / Source ID | Filter / [A] | [ | [ |
AllWISE / | W1 / 33526 | ||
184650.68022907.4 | W2 / 46028 | ||
W3 / 115608 | |||
W4 / 220833 | |||
Spitzer GLIMPSE / | IRAC 3.6 / 35074 | ||
G030.196500.111 | IRAC 4.5 / 44365 | ||
IRAC 5.8 / 56280 | |||
IRAC 8.0 / 75890 | |||
2MASS / | J / 12350 | ||
184650690229072 | H / 16620 | ||
Ks / 21590 | |||
Gaia DR3 / | BP / 5035 | ||
4259072330161989504 | G / 5822 | ||
RP / 7619 | |||
Tycho-2 / | B / 4280 | ||
TYC2 5118-160-1 | V / 5340 | ||
XMM-OM / | UVM2 / 2326 | ||
@@
We used VOSA to fit the deredenned fluxes with a black-body function. The fit parameters are the effective temperature and the dimensionless normalization :
@@
where is the stellar radius and the distance. In Fig.ā7 we show the fluxes of Tableā2 and the result of the fit. We obtain (error correspond to 90% confident level). The stellar flux is and the bolometric luminosity is for a distance of . Using the best-fit value of , we obtain a stellar radius .
To determine the spectral type of BD02 4739, we constructed the HR diagram shown in Fig. 8, where we included the location of the main sequence, sub-giant, and giant stars, corresponding to luminosity classes V, IV and, III, respectively (taken from de Jager & Nieuwenhuijzen 1987). Hence, we suggest that the donor star, i.e. the member that would contribute the most to the IR and optical fluxes, should be a low-mass giant star.
@@ 3.4.2 X-ray spectral analysis
We extracted the X-ray spectra of 4XMM J184650.6022907 from the 3 cameras of each XMM-Newton observation, yielding to a set of 6 spectra. The source extraction region was a circle of Ā in radius and background spectra were extracted from nearby circular regions in the same CCD as the source region. We use the evselect task of SAS with a minimum of 15 counts per bin. We obtained 209 (EPIC-mos1), 245 (EPIC-mos2), and 416 (EPIC-pn) counts for Obs1, and 359 (EPIC-mos1), 391 (EPIC-mos2), and 415 (EPIC-pn) counts for Obs2, in the energy band. The spectra were displayed and analysed with the Xspec software (version 12.14.0), which is included in the HEASoft package.
@@
We visually inspected the 6 spectra and found X-ray emission coming from a soft and featureless continuum, without evidence of spectral lines (Fig. 9). We first fit each observation separately with different emission models in the energy band to look for possible variations of model parameters between Obs1 and Obs2. To explore the non-thermal scenario we use an absorbed power law (). We obtain an unrealistic soft photon index ( and for Obs1 and Obs2, respectively) and a quite poor fit ( and 1.8, for Obs1 and Obs2)888To characterise the goodness of the fit, we report the reduced , i.e. , where are the degrees of freedom.. To explore the thermal nature of the source, we fitted the spectra with a blackbody, a bremsstrahlung, and an optically thin plasma. We obtained poor results for the blackbody ( and 2.0, for Obs1 and Obs2) and bremsstrahlung ( for Obs1 and Obs2), and an acceptable fit for the thin plasma (). For this model, the 90% confidence range for the hydrogen column densities are (Obs1) and (Obs2), showing no significant variation of the absorbing material between observations. For the electron temperature (), we obtain confidence ranges (Obs1) and (Obs2), implying a slightly warmer plasma for Obs2. Metal abundance () is sub-solar and confidence ranges are for Obs1 and for Obs2.
Based on the results of the individual fit of each observation, and in order to increase the fit statistic, we simultaneously fit Obs1 and Obs2 spectra using a thermal plasma model with the same . In Table 3 we show the fitting results for the absorbed thermal plasma, considering two procedures: i) all parameters except normalisation are tie together between Obs1 and Obs2, and ii) only NH is tie together between the observations.
@@
The results of the spectral fitting reported in Table 3 show that the X-ray emission from 4XMM J184650.6022907 is likely produced by a thermal plasma with a temperature and sub-solar abundance. There is a slight increase in the plasma temperature between observations (refer to the right column of the Table), and no significant variations of the metal abundance. The normalisation increases by a factor between observations, i.e. in a time scale of 7 years. This indicates that the increase of the flux is produced by an increase of the amount of emitting plasma.
3.5 IRAS 184430231: a planetary nebula candidate
IRAS 184430231 was first catalogued as a planetary nebula (PN) by Urquhart etĀ al. (2009) as part of the RMS survey. Anderson etĀ al. (2011), in their Green Bank Telescope HiiĀ region survey proposed this source as a āpossibleā PN. Later, it was classified as a PN by Cooper etĀ al. (2013) from its near-infrared spectrum and the absence of millimetre dust continuum emission. Finally, in a near-infrared survey of the inner Galactic plane for Wolf-Rayet stars, Kanarek etĀ al. (2015) also suggested a PN nature for this source.